Determine the density of states in a superconductor.



1
Expert's answer
2021-06-28T13:12:04-0400

The density of states as a function of energy is 


"\frac{dN}{dE}=\frac{V}{2\pi^2}\left(\frac{2m}{h^2}\right)^{\frac{3}{2}}\sqrt{E}"